The temperature coefficient of resistance for carbon at 0oC is -0.0048/oC. What is the significance of the minus sign?
1
Expert's answer
2020-05-05T18:43:27-0400

Temperature coefficient of electrical resistance - a value equal to the relative change in electrical resistance (dR /R) of a portion of an electrical circuit (or specific electrical resistance of a substance) with a change in temperature (dT) per unit:


"\alpha = \frac 1 R \frac {dR}{dT}"


For most metals, the temperature coefficient of resistance is positive: their resistance increases with temperature due to scattering of electrons by phonons.

Negative Temperature Coefficient (NTC) refers to materials that experience a decrease in electrical resistance as their temperature rises. For example, for carbon, resistance decreases with increasing temperature:


"\alpha=-0.0048 \space \frac {1}{\degree C} \space at \space 0\degree C"

A negative temperature coefficient is more often found in semiconductors without impurities, since with increasing temperature the number of electrons in the conduction band increases and the hole concentration accordingly increases. The same temperature dependence of the resistance of solid and nonpolar liquid dielectrics.

NTC materials are used to create inrush current limiters (since they have a higher initial resistance, and when the electric current through the material increases as a result of heating, the electric resistance of the material decreases), temperature sensors and thermistors.